### v2

# A Multi-Level Superoptimizer for Tensor Programs

Abstract.
We introduce Mirage, the first multi-level superoptimizer for tensor programs. A key idea in Mirage is $\mu$Graphs, a uniform representation of tensor programs at the kernel, thread block, and thread levels of the GPU compute hierarchy. $\mu$Graphs enable Mirage to discover novel optimizations that combine algebraic transformations, schedule transformations, and generation of new custom kernels. To navigate the large search space, Mirage introduces a pruning technique based on abstraction that significantly reduces the search space and provides a certain optimality guarantee. To ensure that the optimized $\mu$Graph is equivalent to the input program, Mirage introduces a probabilistic equivalence verification procedure with strong theoretical guarantees. Our evaluation shows that Mirage outperforms existing approaches by 1.1-2.9$\times$ even for DNNs that are widely used and heavily optimized. Mirage is publicly available at https://github.com/mirage-project/mirage.
1. Introduction
Enabling high-performance execution of deep neural networks (DNNs) on GPUs is critical for modern ML applications. Today’s DNN frameworks generally specify DNN computation using tensor programs, which are directed acyclic graphs whose nodes and edges represent tensor algebra operators (e.g., matrix multiplications) and tensors (i.e., $n$-dimensional array) shared between operators.
To optimize an input tensor program, existing frameworks (e.g., PyTorch (pytorch) and TensorFlow (Tensorflow)) use manually designed rules to map the tensor program to expert-written GPU kernels. These approaches generally require extensive engineering efforts to design and implement optimization rules, and may miss some optimization opportunities. To address these challenges, recent work introduced automated approaches to optimizing tensor programs by searching over a comprehensive space of program transformations and applying them based on their performance on target GPUs. These approaches generally fall into two categories.
The first category of work, including Halide (halide), TVM (tvm), and Ansor (ansor), is motivated by the idea of algorithm and schedule separation [^1] In the schedule optimization literature, an algorithm describes what to compute in a kernel and a schedule specifies how to compute the kernel. introduced in Halide and optimizes the schedule of a tensor program while fixing the algorithm. For a given algorithm, these optimizers automatically generate performant kernels by searching for possible strategies to execute the kernel on the target hardware. However, due to the linear algebra nature of DNNs, a tensor program can be represented by a wide spectrum of mathematically equivalent algorithms, and existing schedule-based optimizers only consider kernels whose algorithms are manually specified by users, resulting in missed optimization opportunities.
The second category of work, including TASO (TASO), Grappler (tensorflow_grappler), Tensat (yang2021equality), and PET (wang2021pet), considers algebraic transformations, which exploit mathematical equivalence between different algorithms for a tensor program. Example algebraic transformations include (1) converting a linear algebra operator into another such as a convolution to a matrix multiplication, (2) fusing multiple operators to reduce memory access and kernel overhead, and (3) reorganizing operators based on commutativity, associativity, and distributivity. These optimizers perform algebraic transformations at the algorithm level and require programmers to manually define the set of available kernels and their implementations. They are thus limited by the performance of the provided kernels.
All existing automated optimization approaches, from both categories, still require programmers to manually specify a set of kernels (each defined by a tensor function), and then explore the search space of algebraic or schedule transformations. However, some advanced performance optimizations require coordinated transformations at the kernel, thread block, and thread levels of the GPU compute hierarchy, and involve introducing completely new kernel computations (e.g., a custom kernel that decomposes standard kernels and fuses only some of their computations). Such optimizations are not part of the search space of existing automated methods and must still be implemented manually.
One such example is FlashAttention (dao2023flash) (see §8.2 for details), which optimizes attention (transformers) on GPUs by reordering operators at the algorithm level (algebraic transformations), reorganizing the computation across GPU kernels (yielding new custom kernels), and adapting the parallelization strategy of each kernel to the GPU architecture (schedule transformations). The transformations required for this example cannot be automatically discovered by existing frameworks and must therefore be implemented manually. An implementation of FlashAttention in Triton (tillet2019triton), a widely used tensor program optimizer, includes more than 700 lines of code (triton-flashattention).
We present Mirage, the first multi-level superoptimizer for tensor programs. Mirage is able to automatically discover and verify sophisticated optimizations of tensor programs that require joint optimization of algebraic transformations, schedule transformations, and the discovery of new custom kernels.
A key idea in Mirage is $\mu$Graphs, a hierarchical graph representation that specifies a tensor program at multiple levels of the GPU compute hierarchy. By uniformly treating the kernel, thread block, and thread levels, $\mu$Graphs can capture both algebraic and schedule transformations across these levels. Moreover, optimizing a $\mu$Graph can introduce new custom kernels, which go beyond both algebraic and schedule transformations. For example, Mirage automatically discovers the $\mu$Graphs that represent FlashAttention (dao2023flash) and its inference variant FlashDecoding (tri2023flashdecoding) as well as other $\mu$Graphs that outperform these manually designed kernels by up to 3.5$\times$ for certain use cases. Most of these Mirage-discovered optimizations are outside the search space of existing approaches.
Figure 1. An overview of Mirage.
Figure 1 shows an overview of Mirage. Mirage first splits an input tensor program into subprograms that fall into the restricted Lax fragment. The Lax fragment, formally defined in §5, includes multi-linear operators such as matrix multiplication and convolution, division (which is useful for normalizations), and limited use of exponentiation (which is useful for activations). The partitioning into Lax subprograms reduces the optimization search space while preserving most optimization opportunities; it also enables Mirage’s probabilistic equivalence verifier.
Expression-guided $\mu$Graph generator.
For each Lax subprogram, Mirage’s expression-guided generator uses an exhaustive search to find possible $\mu$Graphs equivalent to it. A key challenge Mirage has to address is its significantly larger search space compared to prior superoptimization techniques for ML. For example, TASO (TASO) and PET (wang2021pet) only search for tensor programs at the kernel level by using a fixed set of pre-defined kernels, while Mirage simultaneously considers superoptimization at the kernel, thread block, and thread levels. To efficiently navigate this significantly larger and more complex search space, Mirage introduces a novel pruning technique based on abstract expressions. This approach greatly reduces the number of $\mu$Graphs Mirage has to consider while providing a theoretical guarantee on the optimality of the discovered $\mu$Graphs (§4.3).
Probabilistic equivalence verifier.
For a $\mu$Graph discovered by Mirage, verifying its functional equivalence with the input program introduces another challenge, since the input and output tensors of a program include up to many millions of elements. A key idea behind Mirage is probabilistic equivalence verification, which performs random tests over finite fields to check equivalence between $\mu$Graphs. While random tests can hardly provide any correctness guarantees for general programs, Mirage relies on a novel theoretical result to show that the restrictions of the Lax fragment ensure that for Lax programs, random tests over finite fields offer strong correctness guarantees. We show that an algorithm for polynomial identity testing (PIT) (schwartz1980fast; zippel1979probabilistic) can be generalized to Lax programs, yielding a randomized algorithm for Lax program equivalence that can be made arbitrarily precise. Mirage uses this random algorithm to (probabilistically) ensure that each optimized program is equivalent to the input program.
$\mu$Graph optimizer.
For each verified $\mu$Graph, Mirage’s $\mu$Graph optimizer maximizes its runtime performance by further considering potential tensor layouts, scheduling operators’ execution orders, and planning memory allocation at all of the kernel, thread block, and thread levels. Finally, Mirage returns an optimized tensor program based on the best discovered $\mu$Graph for each individual Lax subprogram.
Evaluation results.
We evaluate Mirage on a variety of commonly used DNN benchmarks on NVIDIA A100 and H100 GPUs. Even for DNN benchmarks that are widely used and heavily optimized by existing systems such as the group-query attention used in LLMs (grattafiori2024llama3herdmodels), Mirage still outperforms current approaches by 1.1-2.9$\times$, by exploiting subtle custom kernels and optimizations missing in existing systems.
2. Multi-Level Graph Representation
Mirage uses a $\mu$Graph to specify the execution of a tensor program on GPUs. A $\mu$Graph contains hierarchical graphs at multiple levels to represent computation at the kernel, block, and thread levels [^2] For simplicity, we use the term block to refer to a thread block of a CUDA kernel and thread to refer to a single CUDA thread.. This section first describes the GPU hierarchy and uses Figure 4 as a running example to introduce the key components of a $\mu$Graph.
Figure 2. GPU compute and memory hierarchy.
GPU hierarchy.
Figure 2 shows the hierarchy of today’s GPUs. Computations on GPUs are organized as kernels, each of which is a function executed simultaneously on multiple GPU cores in a single-program-multiple-data (SPMD) fashion. A kernel includes a grid of thread blocks, each of which is executed on one GPU streaming multiprocessor and includes multiple threads to perform computation on individual data elements. Each thread is associated with a per-thread register file, and all threads within a thread block can access shared memory to enable collective operations. Finally, all inputs and outputs of a kernel are stored in GPU device memory.
Kernel graph.
Each tensor program corresponds to one kernel graph, where each node represents a kernel running on an entire GPU and each edge is a tensor shared between kernels. All tensors in a kernel graph are stored in GPU device memory since different kernels cannot share data in register file or shared memory. Each node in a kernel graph can be a pre-defined kernel operator supported by existing kernel libraries such as convolution by cuDNN (cudnn) and matrix multiplication by cuBLAS (cublas). In addition, to enable fine-grained inter-kernel optimizations such as kernel fusion, a node in a kernel graph can also be a graph-defined kernel operator, whose semantic and behavior are defined by a lower-level (i.e., block) graph. As an example, the kernel operator in Figure 4b is a graph-defined operator specified by a block graph.
Block graph.
A block graph specifies computation associated with a thread block [^3] In the CUDA programming model, a kernel’s computation is defined as computations for independent thread blocks., where each node denotes a block operator which specifies computation within a block and each edge (blue arrows in Figure 4b) is a tensor shared between block operators. Mirage saves all intermediate tensors within a block graph in GPU shared memory for two considerations. First, GPU shared memory offers much higher bandwidth than device memory, and this design allows Mirage to reduce device memory access by maximally saving intermediate results in shared memory. Second, for tensors whose sizes exceed shared memory capacity and must be stored in device memory, Mirage uses these tensors to split computation into multiple block graphs, each of which only contains tensors in shared memory. This separation does not introduce additional access to device memory.
Each block graph is also associated with a few properties to specify its execution, which we introduce as follows.
(a) $\mbox{\rm\em imap}=\{x\leftrightarrow\mbox{\rm\em column}\}$, $\mbox{\rm\em fmap}=\{i\leftrightarrow\mbox{\rm\em row}\}$
(b) $\mbox{\rm\em imap}=\{x\leftrightarrow\phi,y\leftrightarrow\mbox{\rm\em row}\}$, $\mbox{\rm\em fmap}=\{i\leftrightarrow\mbox{\rm\em column}\}$
Figure 3. Demonstrating how an input tensor is partitioned across blocks and for-loop iterations with imap and fmap.
Grid dimensions.
All blocks within a kernel are organized by a mesh with up to 3 dimensions, identified as $x$, $y$, and $z$. A block graph is associated with up to three grid dimensions that specify the number of blocks along the $x$, $y$, and $z$ dimensions. The block graph in Figure 4b launches $128$ blocks.
First, for each input tensor to a graph-defined kernel operator (e.g., $X$, $G$, and $W$ in the kernel graph in Figure 4b), the associated block graph contains an imap, which specifies how the input tensor is partitioned into sub-tensors for individual blocks. For each grid dimension (i.e., $x$, $y$, or $z$), the imap maps it to (1) a data dimension of the input tensor or (2) a special replica dimension $\phi$. For (1), the mapped data dimension is equally partitioned across blocks along the grid dimension. For (2), the input tensor is replicated across these blocks. As an example, the block graph in Figure 4b takes three inputs — $\overline{X}$, $\overline{G}$, and $\overline{W}$, which represent the input tensors to each block. For $\overline{W}$, its $\mbox{\rm\em imap}=\{x\leftrightarrow d\}$ indicates that the $d$ dimension of tensor $W$ is partitioned into 128 equally sized chunks. As a result, $\overline{W}$ is of shape $[h=1024,d=32]$.
Second, for each output tensor of a block graph (e.g., $\overline{Z}$ in Figure 4b), the block graph includes an omap, which specifies how the outputs of all blocks are concatenated to construct the final output of the kernel operator. In an omap, each grid dimension must map to a data dimension of the output tensor, since different blocks must save to disjoint tensors in device memory. For $\overline{Z}$ of shape $[b=16,d=32]$ in Figure 4b, its $\mbox{\rm\em omap}=\{x\leftrightarrow d\}$ indicates that blocks with the same $x$ index are concatenated along the $d$ dimension, resulting in a tensor $Z$ of shape $[b=16,d=4096]$.
For-loop dimensions.
To fit large input tensors in shared memory and allow cache reuse, a second property associated with each block graph is for-loop dimensions, which collectively specify how many times the block graph is executed to complete a kernel. For example, the block graph in Figure 4b has a for-loop dimension $i=16$, indicating it is executed 16 times to finish the associated graph-defined kernel operator. Correspondingly, each input tensor to a block graph is first sent to an input iterator that loads a part of the tensor (e.g., $\overline{X}$, $\overline{G}$, and $\overline{W}$) from device memory to shared memory. Each input iterator is associated with an fmap to specify which part of the input tensor to load in each iteration. Formally, the fmap maps each for-loop dimension to (1) a data dimension of the input tensor or (2) the replica dimension $\phi$. Similar to the semantic of imap, the input tensor is equally partitioned along that dimension for (1) and replicated for (2). Figure 3 shows how an input matrix is partitioned across blocks and for-loop iterations with different imap and fmap.
In addition, a block graph contains for-loop accumulators (e.g., the two Accum operators in Figure 4b) to accumulate its output across iterations in shared memory. Similar to an input iterator, a for-loop accumulator is also associated with an fmap to specify how the output tensors of different iterations are combined to produce the accumulated results. Specifically, the fmap maps each for-loop dimension to either a data dimension, which results in concatenation of the output along that dimension, or the replica dimension $\phi$, which results in the output being accumulated in shared memory.
Thread graph.
A thread graph further reduces the scope of computation from a block to a single thread. Similar to a block graph, each thread graph is also associated with block dimensions, which specify the organization of threads within the block, and for-loop dimensions, which define the total number of iterations to finish the defined computation. Each thread graph includes input iterators, each of which loads an input tensor (e.g., $\overline{\overline{A}}$ and $\overline{\overline{B}}$ in Figure 4b) from GPU shared memory to register file, and output savers, each of which saves an output tensor from register file back to shared memory (e.g., $\overline{\overline{C}}$). A thread graph is the lowest level graph in a $\mu$Graph and only contains pre-defined thread operators.
Tensor layout.
Each tensor in the kernel, block, or thread graph is associated with a tensor layout (omitted in Figure 4 for simplicity), which specifies how the tensor is linearized in memory. Note that tensor layout only affects the execution performance of a $\mu$Graph and has no impact on its output.
Comparison with prior work.
Prior work separately considers algebraic (TASO; wang2021pet) or schedule transformations (halide; tvm; autohalide), while $\mu$Graphs can represent both in a uniform way. Specifically, the grid and for-loop dimensions and their corresponding mappings (i.e., imap, omap, and fmap) to tensor dimensions constitute a comprehensive search space of possible schedules for graph-defined operators. The hierarchical graphs at the kernel, block, and thread levels allow Mirage to explore algebraic transformations at these levels.
3. Case Study: RMSNorm
(a) Computation graph for RMSNorm and MatMul.
(b) The best $\mu$Graph discovered by Mirage.
Figure 4. Figure 4a is the computation graph for RMSNorm and MatMul. Figure 4b shows the best $\mu$Graph discovered by Mirage for computing RMSNorm and MatMul, which fuses the computation in a single kernel to reduce device memory access and kernel launch overhead, outperforms existing approaches by 1.9$\times$. Numbers in brackets indicate tensor shapes, and numbers in braces show the imap, omap, or fmap for the corresponding operators.
In this section, we use the root mean square layer normalization (RMSNorm) (zhang2019rootmeansquarelayer) as a case study to demonstrate the advantages of the $\mu$Graph representation and Mirage’s superoptimization approach. RMSNorm is a widely used normalization technique by recent large language models such as LLAMA-2-70B (llama2). Formally, RMSNorm takes two tensors $X$ and $G$ as inputs and regularizes the summed inputs of $X$ according to the root mean square static:
(1) $Y_{i}=\frac{X_{i}G_{i}}{\textrm{RMS}(X)},\textrm{RMS}(X)=\sqrt{\frac{1}{d}\sum_{j=1}^{d}X_{j}^{2}},$
where $d$ is the hidden dimension size of $X$.
Figure 5. An overview of the $\mu$Graph generator.
RMSNorm are generally followed by a matrix multiplication (MatMul). Figure 4a shows the computation graph of an RMSNorm and a MatMul operator, where $X$ is the input tensor, and $G$ and $W$ denote two weight tensors. Existing ML compilers generally launch two separate kernels for RMSNorm and MatMul computations, since both of them internally perform reduction across an input dimension, making it challenging to fuse their computations in a single kernel. This approach requires storing intermediate results (i.e., $Y$) in device memory as different kernels cannot share data in shared memory or register files.
Figure 4b shows the best $\mu$Graph automatically discovered by Mirage for computing RMSNorm and MatMul in a single kernel. The computation is fused in a single graph-defined kernel operator to avoid saving intermediate results (i.e., $Y$) in device memory and reduce kernel launch overheads.
We highlight the key differences between the $\mu$Graph discovered by Mirage and the original $\mu$Graph. These differences involve discovering new custom kernels and combining algebraic and schedule transformations, making it infeasible to discover the final $\mu$Graph by separately considering algebraic and schedule transformations. First, Mirage reorders MatMul and the division of RMSNorm by leveraging the commutativity of matrix multiplication and element-wise division (algebraic transformation). Second, Mirage performs the accumulation in the root mean square (i.e., $\sum_{j}X_{j}^{2}$) and the accumulation in the matrix multiplication (i.e., $\sum_{i}X_{i}G_{i}W_{ik}$) in parallel (schedule transformation), avoiding writing the accumulation results to device memory. Next, Mirage instantiates a thread graph to perform a sequence of element-wise operators while maintaining all intermediate results in register files (schedule transformation). Finally, the best discovered $\mu$Graph uses a new custom kernel to fuse the computation of RMSNorm and MatMul, reducing device memory access and kernel launch overheads. This $\mu$Graph outperforms the hand-written kernels in existing systems by 1.9$\times$ and 1.6$\times$ on NVIDIA A100 and H100 GPUs.
4. Expression-Guided $\mu$Graph Generator
Algorithm 1 Mirage’s hybrid $\mu$Graph generation algorithm.
1:A Lax program with a computation graph $G_{\mathsf{ref}}$
2:A set of $\mu$Graphs $\mathcal{S}$
3:$E_{O}\leftarrow\mathrm{AbstractExpression}(G_{\mathsf{ref}})$
4:$\mathcal{S}_{0},\mathcal{S}\leftarrow\varnothing,G\leftarrow\mathrm{Inputs}(G_{\mathsf{ref}})$
5:GenerateKernelOrBlockOps($\mathcal{S}_{0},G$)
6:ConstructThreadGraphs($\mathcal{S}_{0},\mathcal{S}$)
7:
8:function GenerateKernelOrBlockOps($\mathcal{S}_{0},G$)
9:  $\mathcal{S}_{0}\leftarrow\mathcal{S}_{0}\cup\{G\}$
[…]
12:    GenerateNextKernelOperator($G\cup\{o\}$)
[…]
18:  return Invalid
19:function ConstructThreadGraphs($\mathcal{S}_{0},\mathcal{S}$)
20:  $\mathcal{P}\leftarrow$ pre-defined patterns
21:  for all $G\in\mathcal{S}_{0}$ do
22:   $G_{\mathsf{fused}}\leftarrow G$
23:   for all $(G_{i},O_{i})\in\mathcal{P}$ do
24:    for all subgraph $G^{\prime}$ of $G$ matching $G_{i}$ do
25:     Substitute $G^{\prime}$ with $O_{i}$ in $G_{\mathsf{fused}}$
26:   $S\leftarrow S\cup\{G_{\mathsf{fused}}\}$
This section introduces the Mirage $\mu$Graph generator, which automatically discovers potential $\mu$Graphs for an input tensor program. To generate $\mu$Graphs that capture optimizations at all of the kernel, block, and thread levels, Mirage must explore a significantly larger search space than existing superoptimizers that only consider optimizations at the kernel level. Mirage employs two key techniques to address this challenge. First, based on an important observation that optimizations at the kernel and block levels are substantially more critical to performance than optimizations at the thread level since accessing device and shared memory is orders of magnitude more expensive than accessing register file, Mirage’s $\mu$Graph generator employs a hybrid approach, considering all possible graphs up to a certain size at the kernel and block levels, and using a rule-based strategy to construct graphs at the thread level, which reduces the search space while retaining most performance optimizations. Second, to further prune the search space, Mirage introduces a pruning technique based on an abstraction of $\mu$Graphs called abstract expression, which reduces the number of $\mu$Graphs Mirage has to consider while providing a theoretical guarantee on the optimality of the discovered $\mu$Graphs.
4.1. Kernel and Block Graph Generation
[…]
Mirage checks tensor shape (14) and memory usage (15) before adding an operator, ensuring a valid prefix. A prefix passes the memory usage check if: (1) all tensors in the kernel graph can reside in device memory; and (2) all tensors in each block graph can fit in shared memory.
[…]
In addition, Mirage utilizes the abstract expression technique to prune out prefixes that do not satisfy certain constraints, which will be introduced in §4.3.
4.2. Thread Graph Construction
Mirage constructs thread graphs in a way similar to operator fusion, as shown in the third part of Figure 5 and 8-12 in Algorithm 1. Specifically, Mirage has a set of pre-defined patterns $\{(G_{i},O_{i})\}$ where each $G_{i}$ is a graph consisting of block operators and $O_{i}$ is a thread graph. Given a $\mu$Graph, Mirage traverses all its block graphs, replacing any subgraphs that match $G_{i}$ with an operator defined by $O_{i}$.
Table 1. Operators supported by Mirage. The second column shows the levels of graphs supporting the operator (K, B and T stand for kernel, block and thread graphs, respectively). The last column defines the abstract expressions of the outputs of each operator. $E$ is the mapping from tensors to the corresponding abstract expressions.
$\mu$Graph Graph Abstract Expression of Output Tensor
Operator Level
InIter B $\mathrm{E}(\mathrm{InIter}(X))=\mathrm{E}(X)$
OutSaver B $\mathrm{E}(\mathrm{OutSaver}(X))=\mathrm{E}(X)$
Matmul K, B, T $\mathrm{E}(\mathrm{Matmul}(X,Y))=\mathsf{sum}(k,\mathsf{mul}(\mathrm{E}(X),\mathrm{E}(Y)))$1
Sum K, B, T $\mathrm{E}(\mathrm{Sum}(d_{r},k_{r},X))=\mathsf{sum}(k_{r},\mathrm{E}(X))$2
EwAdd K, B, T $\mathrm{E}(\mathrm{EwAdd}(X,Y))=\mathsf{add}(\mathrm{E}(X),\mathrm{E}(Y))$
EwMul K, B, T $\mathrm{E}(\mathrm{EwMul}(X,Y))=\mathsf{mul}(\mathrm{E}(X),\mathrm{E}(Y))$
EwDiv K, B, T $\mathrm{E}(\mathrm{EwDiv}(X,Y))=\mathsf{div}(\mathrm{E}(X),\mathrm{E}(Y))$
EwExp K, B, T $\mathrm{E}(\mathrm{EwExp}(X))=\mathsf{exp}(\mathrm{E}(X))$
Repeat K, B $\mathrm{E}(\mathrm{Repeat}(X))=\mathrm{E}(X)$
Reshape K, B $\mathrm{E}(\mathrm{Reshape}(X))=\mathrm{E}(X)$
Sqr K, B $\mathrm{E}(\mathrm{Sqr}(X))=\mathsf{mul}(\mathrm{E}(X),\mathrm{E}(X))$
[…]
Accum B $\mathrm{E}(\mathrm{Accum}(X,m,i))=\mathsf{sum}(i,\mathrm{E}(X))$ if $m=\phi$ else $\mathrm{E}(X)$ 3
  • 1
    $k$ means the size of the last dimension of $A$, i.e., the reduction dimension. Matmul is performed on the inner most two dimensions and leading dimensions are batched.
  • 2
    Sum along the dimension $d_{r}$ for every $k_{r}$ elements.
  • 3
    Accumulate the results of $i$ for-loop iterations along fmap $m$.
4.3. Pruning via Abstract Expressions
When searching the space of possible $\mu$Graphs, we aim to avoid $\mu$Graph prefixes whose intermediate results cannot contribute to the desired computation. For example, for the input program $X\cdot Z+Y\cdot Z$, we can prune a prefix that computes $X\cdot Y$, but we should not prune one that computes $X+Y$, as $(X+Y)\cdot Z$ is equivalent to the input program. However, how can we determine whether a prefix can contribute to a desired computation while searching for that computation? Below, we develop a pruning technique driven by this intuition that circumvents the “chicken and egg” problem via abstraction. First, we present the abstraction — abstract expressions — and then explain how it is used for pruning. Finally, we offer a theoretical guarantee that under certain conditions, this pruning does not exclude the optimal $\mu$Graph.
Figure 6. Illustration of abstract expressions. The abstract expressions of tensors are annotated on edges. A human-friendly notation is used here: $\mathrm{e}^{a}$ denotes $\mathsf{exp}(a)$, $\sum_{k}a$ denotes $\mathsf{sum}(k,a)$, $a/b$ denotes $\mathsf{div}(a,b)$, and $a*b$ denotes $\mathsf{mul}(a,b)$. The tensors $I_{1}$, $I_{2}$ and $O$ are all $64\times 64$ matrices.
Abstract expressions.
Recall that each edge in a $\mu$Graph corresponds to a tensor-valued function of the input tensors. Intuitively, abstract expressions abstract these functions by ignoring the differences between elements from the same input tensor. Formally, abstract expressions are first-order logic terms over the theory of integers and uninterpreted functions. In a $\mu$Graph, the abstract expression of each edge, denoted by $\mathrm{E}(\cdot)$, is defined as shown in Table 1. During the computation of a $\mu$Graph’s abstract expression, all graph-defined operators are “inlined”. Specifically, the expressions computed for a graph-defined operator’s inputs serve as inputs to the lower-level graph, and the expressions for the outputs of this lower-level graph become the output expressions of the graph-defined operator. Figure 6 shows the abstract expressions for a subgraph of attention.
While abstract expressions capture some information about the function computed at each edge, they also abstract away much of it. For example, if $X$ is a $k\times k$ matrix, summing over the rows and summing over the columns lead to the same abstract expression—$\mathsf{sum}(k,\mathrm{E}(X))$. But keeping $k$ as part of the abstract expression is crucial for effective pruning.
Abstract subexpression and pruning.
We use abstract expressions to prune the search space of $\mu$Graphs by formalizing two relations over abstract expressions: equivalence and abstract subexpression. We prune every $\mu$Graph prefix whose abstract expression is not a subexpression of some abstract expression equivalent to that of the input program. We formalize abstract expressions as uninterpreted functions in first-order logic over the theory of integer arithmetic and uninterpreted functions, and use an SMT solver to reason about them based on two sets of axioms in Table 2: $A_{\text{eq}}$ and $A_{\text{sub}}$. $A_{\text{eq}}$ axiomatizes equivalence between abstract expressions. As will become clear below, these axioms need not be sound—it is not necessary that $\mu$Graphs whose abstract expressions are equivalent are actually equivalent. (As mentioned earlier, non-equivalent $\mu$Graphs can have identical abstract expressions.) $A_{\text{sub}}$ axiomatizes the subexpression relation between abstract expressions. The key property of this relation is that whenever a $\mu$Graph $G_{1}$ is a prefix of $G_{2}$, meaning $G_{2}$ can be obtained from $G_{1}$ by adding more operators, then $\mathrm{E}(G_{1})$ is an abstract subexpression of $\mathrm{E}(G_{2})$, i.e., $A_{\text{sub}}\models\mathsf{subexpr}(\mathrm{E}(G_{1}),\mathrm{E}(G_{2}))$, where $\models$ denotes entailment modulo the theory of integer arithmetic and uninterpreted functions.
During the search, Algorithm 1 first computes the abstract expression of the input Lax program, $E_{O}$, and prunes away any $\mu$Graph prefix $G$ if $A_{\text{eq}}\cup A_{\text{sub}}\not\models\mathsf{subexpr}(\mathrm{E}(G),E_{O})$. That is, we prune a graph if its abstract expression is not a subexpression of $E_{O}$. The check is performed by invoking an SMT solver (Z3 (z3)). As an optimization, check results are cached and reused, since during the search Mirage may encounter multiple $\mu$Graphs with identical abstract expressions and SMT queries are relatively expensive.
Table 2. Axiomatization of abstract expressions used for pruning. Mirage checks whether an abstract expression $E_{1}$ is a subexpression of $E_{2}$ by querying an SMT solver to check if $\mathsf{subexpr}(E_{1},E_{2})$ is entailed by these axioms. All variables are universally quantified.
Abstract Expression Property
Equivalence Axioms $A_{\text{eq}}$ Subexpression Axioms $A_{\text{sub}}$
$\mathsf{add}(x,y)=\mathsf{add}(y,x)$ $\mathsf{subexpr}(x,\mathsf{add}(x,y))$
$\mathsf{mul}(x,y)=\mathsf{mul}(y,x)$ $\mathsf{subexpr}(x,\mathsf{mul}(x,y))$
$\mathsf{add}(x,\mathsf{add}(y,z))=\mathsf{add}(\mathsf{add}(x,y),z)$ $\mathsf{subexpr}(x,\mathsf{div}(x,y))$
$\mathsf{mul}(x,\mathsf{mul}(y,z))=\mathsf{mul}(\mathsf{mul}(x,y),z)$ $\mathsf{subexpr}(y,\mathsf{div}(x,y))$
$\mathsf{add}(\mathsf{mul}(x,z),\mathsf{mul}(y,z))=\mathsf{mul}(\mathsf{add}(x,y),z)$ $\mathsf{subexpr}(x,\mathsf{exp}(x))$
$\mathsf{add}(\mathsf{div}(x,z),\mathsf{div}(y,z))=\mathsf{div}(\mathsf{add}(x,y),z)$ $\mathsf{subexpr}(x,\mathsf{sum}(i,x))$
[…]
$x=\mathsf{sum}(1,x)$ $\mathsf{subexpr}(x,x)$
$\mathsf{sum}(i,\mathsf{sum}(j,x))=\mathsf{sum}(i*j,x)$ $\mathsf{subexpr}(x,y)\wedge\mathsf{subexpr}(y,z)$
$\mathsf{sum}(i,\mathsf{add}(x,y))=\mathsf{add}(\mathsf{sum}(i,x),\mathsf{sum}(i,y))$ $\ \to\mathsf{subexpr}(x,z)$
$\mathsf{sum}(i,\mathsf{mul}(x,y))=\mathsf{mul}(\mathsf{sum}(i,x),y)$
$\mathsf{sum}(i,\mathsf{div}(x,y))=\mathsf{div}(\mathsf{sum}(i,x),y)$
Theoretical guarantee and the pruning-optimality tradeoff.
Intuitively, our pruning would keep any prefix that can lead to a $\mu$Graph whose abstract expression is equivalent (according to $A_{\text{eq}}$) to that of the input Lax program. Formally:
Theorem 1 (Pruning via Abstract Expressions).
For an input $\mu$Graph $G_{0}$, and a $\mu$Graph $G$ equivalent to $G_{0}$, if $A_{\text{eq}}\models E(G_{0})=E(G)$ then $G$ will be generated by Algorithm 1.
Proof.
By Tables 1 and 2, we show that for any operator op, if $Y=\mbox{\rm\em op}(X_{1},\ldots,X_{n})$ then $A_{\text{sub}}\models\mathsf{subexpr}(\mathrm{E}(X_{i}),\mathrm{E}(Y))$ (for $1\leq i\leq n)$. That is, an input to an operator is always an abstract subexpression of its output. By the reflexivity and transitivity axioms included in $A_{\text{sub}}$, it follows for any $G^{\prime}$ that is a prefix of $G$, $A_{\text{sub}}\models\mathsf{subexpr}(\mathrm{E}(G^{\prime}),\mathrm{E}(G))$. Together with the assumption that $A_{\text{eq}}\models E(G_{0})=E(G)$, it follows that $A_{\text{eq}}\cup A_{\text{sub}}\models\mathsf{subexpr}(\mathrm{E}(G^{\prime}),\mathrm{E}(G_{0}))$. Thus, any prefix of $G$ will not be pruned, and $G$ will be generated by Mirage. ∎
The theorem highlights the role of abstract expressions in solving the “chicken and egg” problem outlined above. To decide if a prefix $\mu$Graph is useful, we reason about whether it is a prefix of a useful computation in the abstract. The choice of the abstraction and of the axioms $A_{\text{eq}}$ represents a tradeoff between optimality and pruning. As Theorem 1 shows, we are only guaranteed to find the optimal $\mu$Graph such that $A_{\text{eq}}$ imply equivalence of its abstract expression to that of the input program. The stronger the axioms, the theorem covers more $\mu$Graphs, but we also get less pruning because more prefixes would pass the subexpression test. In particular, note that $A_{\text{eq}}$ does not consider cancellation (e.g., $\mathsf{div}(\mathsf{mul}(x,y),y)=y$). As a result, Mirage may miss some equivalent $\mu$Graphs. But including an axiom for cancellation of division and multiplication would make everything a subexpression of everything, therefore nulling the desired pruning. As our evaluation shows, our choice of $A_{\text{eq}}$ yields a good balance between pruning and optimality.
5. Probabilistic Equivalence Verifier
Mirage’s probabilistic equivalence verifier checks if a candidate $\mu$Graph is equivalent to the desired Lax subprogram. The key idea is to evaluate both on random inputs in two finite fields. Using finite fields instead of floating point numbers not only avoids floating point errors, but also leads to a strong theoretical guarantee: the probability of accepting a non-equivalent $\mu$Graph can be made arbitrarily low.
For general programs, random tests can hardly provide any correctness guarantees. However, we show that for Lax programs (formally defined below), random testing provides a probabilistic correctness guarantee, and repeated tests can reduce the error probability to an arbitrarily small threshold.
Prior work (wang2021pet) has applied a similar technique to check equivalence between tensor programs that only contain linear operators (e.g., matrix multiplication, convolution). We develop a random testing technique that also supports division and exponentiation, which are needed for many DNN optimizations (e.g., the RMSNorm example in §3).
Mirage verifies equivalence between Lax $\mu$Graphs (linear, division, and an exponential) defined below. We introduce the main theoretical results in §5.1 and present Mirage’s verification methodology in §5.2.
Definition 1 (Lax $\mu$Graph).
A $\mu$Graph $G$ is a Lax $\mu$Graph if (1) $G$ contains only multi-linear operators [^4] Operator op with $n$ inputs is multi-linear if op is linear to all inputs $I_{k}$:
(1) $\forall X,Y.\mbox{\rm\em op}(I_{1},...,I_{k-1},X,I_{k+1},...,I_{n})+\mbox{\rm\em op}(I_{1},...,I_{k-1},Y,I_{k+1},...,I_{n})=\mbox{\rm\em op}(I_{1},...,I_{k-1},X+Y,I_{k+1},...,I_{n})$, and
(2) $\alpha\cdot\mbox{\rm\em op}(I_{1},...,I_{k-1},X,I_{k+1},...,I_{n})=\mbox{\rm\em op}(I_{1},...,I_{k-1},\alpha\cdot X,I_{k+1},...,I_{n}).$
, division, and exponentiation, and (2) every path from an input to an output in $G$ includes at most one exponentiation.
5.1. Theoretical Foundations
Without loss of generality, we assume a Lax $\mu$Graph $G$ takes $n$ input tensors and produces one output tensor. Our theoretical results can directly generalize to Lax $\mu$Graph with multiple outputs. Since each Lax $\mu$Graph includes linear operators, divisions, and at most one exponentiation along each path, the computation for each entry of the output tensor can be expressed in the following form (by using standard identities such as $\frac{\frac{a}{b}}{\frac{c}{d}}=\frac{ad}{bc}$, $\frac{a}{b}+\frac{c}{d}=\frac{ad+bc}{bd}$, $e^{x}e^{y}=e^{x+y}$):
(2) $\sum_{i=1}^{k}\frac{f_{i}}{g_{i}}e^{\left({\frac{h_{i}}{u_{i}}}\right)},$
where $f_{i}$, $g_{i}$, $h_{i}$, and $u_{i}$ ($1\leq i\leq k$) are polynomials over the entries of the input tensors.
The main theoretical result that underpins our randomized equivalence verification is the following theorem, which extends polynomial identity testing (PIT) (schwartz1980fast; zippel1979probabilistic) on finite fields to Lax $\mu$Graphs. Note that the difference of two Lax $\mu$Graphs is also of the form of Equation 2. Therefore, identity testing of two Lax $\mu$Graphs can be done by testing if an expression of that form is zero. Because of the use of exponentiation, we use two finite fields instead of one. [^5] We use two primes $p$ and $q$ for polynomial identity testing (schwartz1980fast; zippel1979probabilistic) outside and inside the exponents, respectively. The condition $q$ divides $p-1$ is to ensure the existence of $q$-th roots of unity in $\mathbb{Z}_{p}$.
Theorem 2.
Let $\mathbb{Z}_{p},\mathbb{Z}_{q}$ be the finite field of integers modulo $p$ and $q$, respectively, where $p,q$ are primes such that $q$ divides $p-1$. Let $d,k\in\mathbb{N}$ s.t. $dk^{4}<q$. Let $f_{1},\ldots,f_{k},g_{1},\ldots,g_{k}:\mathbb{Z}_{p}^{N}\to\mathbb{Z}_{p}$ be polynomials over $\mathbb{Z}_{p}$ and $h_{1},\ldots,h_{k},u_{1},\ldots,u_{k}:\mathbb{Z}_{q}^{M}\to\mathbb{Z}_{q}$ be polynomials over $\mathbb{Z}_{q}$, where the degrees of all polynomials are at most $d$. Let $\mathcal{G}$ be the set of $q$-th roots of unity in $\mathbb{Z}_{p}$. If $h_{i}/u_{i}\not\equiv h_{j}/u_{j}$ for all $i\neq j$, and $f_{i},g_{i},h_{i}\not\equiv 0$ for all $i$, then
$\Pr_{\mathbf{x}\leftarrow\mathbb{Z}_{p}^{N},\mathbf{y}\leftarrow\mathbb{Z}_{q}^{M},\omega\leftarrow\mathcal{G}}\left[\sum_{i=1}^{k}\frac{f_{i}(\mathbf{x})}{g_{i}(\mathbf{x})}\omega^{\frac{h_{i}(\mathbf{y})}{u_{i}(\mathbf{y})}}=0\wedge\mathcal{E}\right]\leq 1-\frac{1}{k}+o\left(\frac{1}{k}\right),$
where $\mathcal{E}$ is the event that $g_{i}(\mathbf{x}),u_{i}(\mathbf{y})\neq 0$ for all $i$.
5.2. Random Tests over Finite Fields
Table 3. Arithmetic operations for random testing. Mirage selects two prime numbers $p$ and $q$ such that $q$ divides $p-1$. $x_{p}$ and $x_{q}$ are values from $\mathbb{Z}_{p}$ and $\mathbb{Z}_{q}$, respectively.
Opt. Opd. 1 Opd. 2 Output
Add. $(x_{p},x_{q})$ $(y_{p},y_{q})$ $\big{(}x_{p}+y_{p}(\bmod{p}),x_{q}+y_{q}(\bmod{q})$)
Sub. $(x_{p},x_{q})$ $(y_{p},y_{q})$ $\big{(}x_{p}-y_{p}(\bmod{p}),x_{q}-y_{q}(\bmod{q})$)
Mul. $(x_{p},x_{q})$ $(y_{p},y_{q})$ $\big{(}x_{p}y_{p}(\bmod{p}),x_{q}y_{q}(\bmod{q})\big{)}$
Div. $(x_{p},x_{q})$ $(y_{p},y_{q})$ $\big{(}x_{p}y_{p}^{-1}(\bmod{p}),x_{q}y_{q}^{-1}(\bmod{q})\big{)}$
Exp. $(x_{p},x_{q})$ $\big{(}\omega^{x_{q}}(\bmod{p}),-\big{)}$
Mirage leverages Theorem 2 to probabilistically verify the equivalence of two $\mu$Graphs by performing random testing over the finite fields $\mathbb{Z}_{p}$ and $\mathbb{Z}_{q}$ as defined in Theorem 2. To check the equivalence of two $\mu$Graphs, Mirage first generates the input tensors, where each entry is uniformly sampled from $\mathbb{Z}_{p}\times\mathbb{Z}_{q}$. Mirage also samples $\omega$, which is used for exponentiation, uniformly from all the $q$-roots of unity in $\mathbb{Z}_{p}$. Mirage then evaluates the two $\mu$Graphs over the finite fields for the same inputs using the operations defined in Table 3. As explained in §5.1, $\mathbb{Z}_{p}$ and $\mathbb{Z}_{q}$ are used outside and inside the exponent, respectively. All operations except exponentiation are implemented via modular arithmetic on $\mathbb{Z}_{p}$ and $\mathbb{Z}_{q}$ separately. Exponentiation uses the value $x_{q}$ from $\mathbb{Z}_{q}$ and transforms it to a value in $\mathbb{Z}_{p}$ by computing $\omega^{x_{q}}(\bmod p)$. Note that in a Lax $\mu$Graph, exponentiation will be performed at most once along each path. Finally, Mirage checks whether the two $\mu$Graphs produce identical outputs. This process is repeated multiple times, and the two $\mu$Graphs pass the equivalence test if they pass all random tests. The following theorem, which follows from Theorem 2, shows that this process can yield an arbitrarily low error rate.
Theorem 3.
Equivalent $\mu$Graphs always pass $\mu$Graph verification. For two non-equivalent $\mu$Graphs and a given probability threshold $0<\delta\leq 1$, the $\mu$Graphs pass all $\Omega(k\cdot\ln\frac{1}{\delta})$ random tests with probability at most $\delta$.
6. $\mu$Graph Optimizer
For each verified $\mu$Graph, Mirage’s $\mu$Graph optimizer maximizes its performance by further performing layout optimizations, operator scheduling, and memory planning, as shown in Figure 1. Mirage defers these $\mu$Graph optimizations after verification for two reasons. First, these optimizations do not affect the correctness of the generated $\mu$Graphs; omitting them when generating $\mu$Graphs reduces the search space Mirage must consider, since $\mu$Graphs with the same graph topology but different choices of tensor layouts, operator orders, or memory allocation plans are considered identical by the $\mu$Graph generator. Second, considering these optimizations after verification also reduces the search space for these optimizations, since the $\mu$Graph optimizer only needs to optimize $\mu$Graphs that are functionally equivalent to the input.
Tensor layouts.
The $\mu$Graph optimizer considers possible data layouts for all intermediate tensors at the kernel, block, and thread levels and chooses the best combinations to maximize performance. We formulate layout selection as a constrained optimization problem and use an integer linear programming (ILP) algorithm to solve it optimally. Specifically, for each tensor $t$ and each possible layout $l$ for $t$, we introduce a boolean variable $B_{t,l}$ to indicate whether tensor $t$ uses layout $l$. Operators at the kernel, block, and thread levels may impose various constraints on tensor layouts. For example, to use kernels in the cuBLAS library (cublas) to perform matrix multiplication, the innermost dimension of the two input tensors must be among the last two dimensions. These restrictions are converted into a series of linear constraints on $B_{t,l}$. Different tensor layouts may lead to varying performance. For example, some input tensor layouts support bulk copies from device to shared memory, while others do not. Mirage introduces a cost function to model the performance of each operator under different layout choices. Mirage uses an off-the-shelf ILP solver (i.e., Z3 (z3)) to find an optimal layout strategy that satisfies all layout constraints while minimizing cost.
Operator scheduling.
In a $\mu$Graph, there are multiple topological orders to execute operators, and different orders may yield different performance. For a given input $\mu$Graph, the $\mu$Graph optimizer identifies an efficient strategy to schedule operators by minimizing thread-level synchronization within each thread block (i.e., __syncthreads() in CUDA). To achieve this goal, Mirage labels each node with a depth, defined as the length of the longest path from any input operator to that node. Mirage uses a dynamic programming algorithm to compute the depth of each node and schedules all operators in ascending order of their depths. This approach minimizes the number of thread-level synchronizations required in the generated CUDA kernel, as Mirage only needs to insert synchronization points between operators of different depths.
Memory planning.
A third class of post-verification optimizations is memory planning, which determines the memory offsets for all intermediate tensors at the kernel, block, and thread levels. Mirage formulates memory planning as a dynamic storage allocation task and exhaustively enumerates all allocation plans to discover an optimal strategy.
7. Implementation
Table 4. DNN benchmarks used in our evaluation.
Name Description Base Architecture
GQA Group-query attention LLaMA-3-70B (grattafiori2024llama3herdmodels)
QKNorm QK normalization with attention Chameleon-7B (chameleonteam2024chameleon)
RMSNorm RMS normalization with linear LLaMA-2-7B (llama2)
LoRA Low-rank adaptation GPT-3-7B-LoRA (llama-7b-lora)
GatedMLP Gated multi-layer perceptron Falcon-7B (falcon40b)
nTrans Normalized Tarnsformer nGPT-1B (loshchilov2024ngpt)
Mirage is implemented in 30K lines of code in C++, CUDA, and Python. Kernel operators are implemented with the cuDNN and cuBLAS libraries (cudnn; cublas), and block and thread operators are implemented using cuTLASS (cutlass) and CUDA functions. For each input tensor program, Mirage automatically generates and verifies potential $\mu$Graphs. For each verified $\mu$Graph, Mirage produces CUDA source code for all custom kernels of the $\mu$Graph and compiles the code into binary using the CUDA compiler. This approach enables just-in-time (JIT) compilation and deployment for general tensor programs, and the generated kernels can be directly integrated into a PyTorch program with a few lines of code changes. The SMT and ILP solvers of Mirage are implemented using Z3 4.12.6 (z3).
Our implementation supports the operators listed in Table 1. Mirage can be extended to include new operators, such as variants of convolution or matrix multiplication, at the kernel, block, and/or thread levels. To support a new linear operator, Mirage requires (1) a float-pointing implementation of the operator at the kernel, block, and/or thread levels, which is used by the $\mu$Graph optimizer to generate CUDA kernels, (2) an implementation of the operator over modular arithmetic (see §5), and (3) an extension to the abstract expression axioms $A_{\text{eq}}$ and $A_{\text{sub}}$ for it (see §4.3).
To utilize Theorems 2 and 3, random tests should be performed with sufficiently large prime numbers $p$ and $q$ and should be iterated multiple times. Our current implementation uses the largest values of $p$ and $q$ whose product fits in 16-bit integers (i.e., $p=227,q=113$) to perform the random testing on GPUs, leveraging Mirage’s GPU optimizations such as maintaining intermediate results in GPU shared memory, which allows Mirage to accelerate its search procedure. We also use a single random test without iterating it and compare all elements of the output tensors. We note that this equivalence verification procedure does not introduce false negatives. While false positives may be introduced, we have not observed any in practice. For these reasons, we consider this procedure to be sufficient for the search process and plan to add an additional verification step that provides the theoretical guarantees only for the best $\mu$Graph at the end of the optimization process.
Equivalence verification for non-Lax programs.
While Mirage can generate $\mu$Graphs for arbitrary input tensor programs, the probabilistic equivalence verifier is restricted to Lax programs and does not support certain DNN operators such as ReLU (relu). As an alternative, we have developed a solver-based verifier for arbitrary tensor programs. The verifier relies on user-provided mathematical properties of individual operators (e.g., linearity, associativity, commutativity, and distributivity) defined in first-order logic and uses these properties to verify equivalence using an automated theorem prover. Compared to the probabilistic equivalence verifier, the solver-based verifier supports more general programs, while requiring additional manual effort to specify the properties for each new operator. A detailed discussion of the solver-based verifier is beyond the scope of this paper.
8. Evaluation
8.1. Experimental Setup
Figure 7. Comparing Mirage with existing systems for 6 benchmarks on an A100 and an H100 GPU. The performance of all systems are normalized by Mirage (higher is better). Numbers above the Mirage bars show the speedup over the best baselines.
Since Mirage is a superoptimizer for Lax programs, we focus our evaluation on various DNN benchmarks commonly used by existing DNNs, each of which is a Lax program. These benchmarks provide the most fine-grained way to compare the performance of Mirage and existing systems. Table 4 shows the six benchmarks in our evaluation. GQA, RMSNorm, and GatedMLP are the main building blocks of large language models (LLMs). QKNorm introduces query-key normalization before attention to enhance model convergence (chameleonteam2024chameleon). LoRA enables low-rank adaptation for fine-tuning a DNN on different tasks. In addition, we also evaluate how Mirage-generated kernels can improve the end-to-end performance of DNNs, including Chameleon (chameleonteam2024chameleon), nGPT (loshchilov2024ngpt), LLaMA-3 (grattafiori2024llama3herdmodels), and LoRA (hu2021lora).
The experiments were performed on a NVIDIA 40GB A100 and a NVIDIA 40GB H100 GPU. All our benchmarks fit on a single GPU except GQA (used for LLaMA-2-70B), which is generally parallelized across four GPUs using tensor model parallelism (Megatron). Therefore, we evaluate GQA under the tensor model parallelism strategy (i.e., the 8 key-value heads are equally partitioned across four GPUs). Since the performance of Mirage and all baselines depends only on the shapes of the input tensors, we repeat each experiment 1,000 times using random inputs and report the average run time.
[…]
Unless otherwise stated, Mirage considers up to 5 operators in the kernel graph and up to 11 operators in each block graph.
8.2. Benchmark Results
Figure 7 compares the performance of Mirage with existing tensor program optimizers on six DNN benchmarks on NVIDIA A100 and H100 GPUs. All systems use half-precision floating points to serve these DNN benchmarks. TASO (TASO) and PET (wang2021pet) are DNN superoptimizers that automatically generate graph optimizations at the kernel level. PyTorch (pytorch) uses the highly engineered cuDNN and cuBLAS libraries (cublas; cudnn) to perform DNN operators on GPUs. TensorRT and its LLM variant TensorRT-LLM include a set of manually designed and highly optimized kernels for common tensor operators such as attention (tensorrt). FlashAttention and its inference variant FlashDecoding are manually designed and implemented kernels for fast attention computation (dao2023flash; hong2024flashdecoding). Finally, Triton (tillet2019triton) is a schedule-based optimizer to generate high-performance kernels and has been deployed in existing systems, achieving better performance than other schedule-based optimizers.
Compared to the best existing approaches, Mirage improves the performance of these benchmarks by 1.1-2.9$\times$ by combing algebraic transformations, schedule transformations, and the generation of new custom kernels. §3 shows the best discovered $\mu$Graphs for RMSNorm. Next, we present a case study for the other benchmarks.
(a) The kernel graph for QKNorm and attention in existing systems.
(b) The best $\mu$Graph discovered by Mirage for QKNorm and attention.
Figure 8. Comparing the $\mu$Graphs used by existing optimizers and Mirage for QKNorm and attention.
GQA.
Group-query attention is the backbone of LLMs and has been heavily optimized by existing frameworks. For example, FlashAttention and FlashDecoding are expert-designed attention kernels and have been adopted in existing LLM inference systems (dao2023flash). Mirage discovers existing expert-designed kernels as well as other $\mu$Graphs that outperform them by up to 2.2$\times$. The speedup is achieved by two additional optimizations on top of existing hand-written kernels. First, current approaches rely on heuristics to determine the grid dimensions for GQA, which are suboptimal in certain scenarios. For example, TensorRT-LLM launches the attention kernel with grid dimensions of (8, 2, 1) and (8, 2, 8) for GQA when the batch sizes are 1 and 8, respectively. Both configurations cannot fully utilize all A100 and H100 GPUs, which have 108 and 132 SMs respectively. On the other hand, Mirage automatically searches for the best grid dimensions for each $\mu$Graph, resulting in fully utilized SMs.
Second, existing approaches use fixed tensor dimensions to parallelize GQA across thread blocks. For example, FlashAttention (dao2023flash) parallelizes attention across blocks using the sample, head, and query sequence dimensions, while FlashDecoding and TensorRT-LLM leverage the sample, head, and key-value sequence dimensions. Both strategies are efficient for conventional multi-head attention with many attention heads but suboptimal for GQA with limited attention heads. In contrast, Mirage automatically selects the most efficient dimensions among sample, KV heads, query sequence, and key-value sequence, and uses different $\mu$Graphs for different attention scenarios, reducing device memory access by up to 7$\times$ compared to the heuristics used in existing systems. Implementing Mirage’s $\mu$Graphs in existing systems is possible but requires extensive engineering effort to support different kernels for different scenarios, while Mirage automatically generates them and verify their correctness.
QKNorm.
To reduce model divergence, several recent DNNs introduce query-key normalization (QKNorm) into the Transformer architecture (chameleonteam2024chameleon). QKNorm applies layer normalization to the query and key vectors before attention, as shown in Figure 8a. These additional normalization layers are not yet supported by existing attention implementations (e.g., FlashAttention and TensorRT-LLM) and require launching separate kernels for normalization and attention.
Mirage automatically discovers a $\mu$Graph that integrates QKNorm and attention computation into a custom kernel, as shown in Figure 8b. The $\mu$Graph reorganizes the attention computation to enable fusion with the two layer normalizations, which avoids writing intermediate results to GPU device memory and reduces the kernel execution time by 1.1-1.4$\times$.
(a) The kernel graph for LoRA in existing systems.
(b) The best $\mu$Graph discovered by Mirage for LoRA.
Figure 9. Comparing the tensor programs used by existing optimizers and by Mirage for LoRA: $O=W\times X+B\times A\times X$. Note that both matrices $A$ and $B$ are low-rank.
LoRA.
Low-rank adaptation (LoRA) introduces a pair of low-rank adapters to the linear operators of a pre-trained DNN to improve its predictive performance for downstream tasks. Existing tensor program optimizers launch separate kernels for the original linear operator and the two new linear operators in LoRA (Figure 9a), which introduces high kernel launch overheads since the LoRA operators involve very low computational costs. Figure 9b shows the best $\mu$Graph discovered by Mirage for LoRA, which fuses the three Matmuls and the subsequent Add into a single kernel. Mirage reorganizes the computation into two thread-block level Matmuls by leveraging the following algebraic transformation: $W\times X+B\times A\times X=(W\|B)\times\big{(}X\|(A\times X)\big{)}$. The Concats in Figure 9b do not involve any computation and are performed by updating tensor offsets in GPU shared memory. This $\mu$Graph reduces the execution cost of LoRA by 1.7-1.8$\times$.
(a) The kernel graph for GatedMLP.
(b) The best $\mu$Graph discovered by Mirage for GatedMLP.
Figure 10. Comparing the $\mu$Graphs used by existing optimizers and Mirage for GatedMLP.
GatedMLP.
Gated multi-layer perceptron is commonly used in DNNs to capture non-linear representations. We use the GatedMLP configuration introduced in Falcon-7B (falcon40b), whose kernel graph is shown in Figure 10a. Existing tensor program optimizers generally fuse the two Matmuls in a single kernel to reduce GPU device memory access since the input tensor $X$ only needs to be loaded once. However, this approach still requires launching multiple kernels and storing intermediate results (i.e., the output of the two Matmuls) in device memory as the SiLU activation and elementwise multiplication are not fused with the Matmuls.
In contrast, the best $\mu$Graph discovered by Mirage (Figure 10b) performs the two Matmuls in parallel in the same block graph and fuses the remaining computation (i.e., SiLU and Mul) as a post-processing step into the block graph. This approach yields 1.4-1.5$\times$ speedups on A100 GPUs and 2.7-2.9$\times$ speedups on H100 GPUs.
nTrans.
To accelerate model training, nGPT introduces normalized Transformer, which normalizes all intermediate results in Transformer. Formally, the computation is defined as $y=\texttt{Norm}(x+\alpha(\texttt{Norm}(h-x)))$, where Norm is a normalization layer, and $x$, $h$, and $\alpha$ are the input tensors. Existing systems launch three separate kernels for nTrans since it interleaves normalization and elementwise addition and multiplication. Mirage automatically discovers a $\mu$Graph that fuses the computation into a single kernel and stores all intermediate results in the GPU shared memory, reducing the kernel execution time by 1.8-2.0$\times$ compared to the best existing systems.
Figure 11. Comparing the end-to-end inference performance of PyTorch and PyTorch with Mirage-generated kernels.
8.3. End-to-end Results
In addition to the microbenchmark performance, we also evaluate how Mirage-generated kernels reduce the end-to-end latency of commonly used DNNs. Mirage supports just-in-time compilation and deployment, and its generated kernels can be directly integrated in PyTorch programs. We compare PyTorch with its native handwritten CUDA kernels and PyTorch with Mirage-generated kernels on four DNN models. Figure 11 shows the results. Mirage reduces the end-to-end latency of these models by 1.1-1.9$\times$ by automatically generating highly optimized kernels. Note that the improvement is achieved by a few lines of code changes to the PyTorch programs.
8.4. Optimization Time and Ablation Study
Table 5. Ablation study on Mirage’s techniques to accelerate $\mu$Graph generation. We evaluate the impact of multi-threading and abstract expressions on search time for RMSNorm.
Max # Ops in Mirage Mirage w/o Mirage w/o
a Block Graph Multithreading Abstract Expression
$5$ $11$ sec $58$ sec $768$ sec
$6$ $16$ sec $93$ sec $19934$ sec
$7$ $22$ sec $150$ sec $>10$ h
$8$ $24$ sec $152$ sec $>10$ h
$9$ $26$ sec $166$ sec $>10$ h
$10$ $26$ sec $166$ sec $>10$ h
$11$ $28$ sec $183$ sec $>10$ h
Generally, Mirage takes up to 4 hours to optimize a Lax program. Note that this optimization is a one-time cost before deployment on the target hardware. This section provides more detailed results and an ablation study for Mirage’s search procedure. We evaluate how our techniques allow Mirage to explore large $\mu$Graphs while maintaining low search time. We focus on two techniques: pruning via abstract expressions (§4.3) and multi-threading. Table 5 reports the search times for RMSNorm as we adjust the maximum number of operators in a block graph. Multi-threading significantly reduces the search time, while pruning via abstract expressions is crucial for the scalability of Mirage. Specifically, the pruning techniques allow Mirage to explore $\mu$Graphs whose block graph can each have at most 11 operators, while disabling abstract expressions restricts Mirage to consider up to 6 operators in a block graph in order to finish the search in 10 hours. Note that Mirage needs to consider 11 operators in a block graph in order to discover the $\mu$Graph shown in Figure 4 for RMSNorm.
9. Related Work
Manually-designed kernels.
Many existing frameworks, such as TensorFlow XLA (tensorflow_xla; Tensorflow), PyTorch (pytorch), and TensorRT (tensorrt), rely on GPU experts to manually design and implement kernels for ML operators. Recently, significant engineering effort has been dedicated to manually designing, implementing, and optimizing GPU kernels for commonly used DNNs (known as foundation models (bommasani2022opportunities)). For example, to optimize attention computation (transformers), recent work has introduced various kernels based on FlashAttention (dao2023flash; tri2023flashdecoding; hong2024flashdecoding; fastertransformer). Due to the increasing complexity of modern GPU architectures (e.g., tensor cores in A100s (Markidis2018tensorcores) and thread block clusters in H100s (nvidia-h100)), manually designed kernels may miss subtle optimizations that are hard to discover manually.
Superoptimization-based approaches.
Superoptimization was originally introduced to find optimal code for instruction sequences (massalin; stoke; peephole). Recent work has applied superoptimization techniques to tensor programs (TASO; wang2021pet; zheng2023einnet; tensat). All these attempts only consider algebraic transformations at the kernel level and cannot discover sophisticated optimizations that require jointly considering algebraic and schedule transformations at all of the kernel, block, and thread levels. Our evaluation shows that Mirage largely outperforms existing DNN superoptimizers, demonstrating the significance of multi-level joint optimizations.
Schedule-based approaches.
Recent work has introduced ML compilers that automatically optimize the schedule of kernel executions on GPUs. For example, TVM (tvm; tvm_auto_tuner), Ansor (ansor), and Triton (tillet2019triton), among others (zheng2020flextensor; hagedorn2023graphene; feng2022tensorir) are based on the idea of algorithm-schedule separation introduced in Halide and search for an optimized schedule to execute a given algorithm on GPUs. Schedule-based approaches rely on users to explicitly specify the algorithm for each kernel and their performance is limited to the provided algorithms.
Multi-level graph representations.
Welder (shi2023welder) and ASPEN (park2023aspen) introduce multi-level tile graphs that bear some similarity to Mirage’s $\mu$Graphs, as both representations follow the GPU memory hierarchy. However, prior work focuses on scheduling transformations, while Mirage covers these optimizations and also considers algebraic transformations and discovery of new custom kernels. Most optimizations shown in the paper are outside the scope of prior work.
10. Conclusion
This paper proposes Mirage, the first multi-level superoptimizer for tensor programs. Mirage introduces a hierarchy graph representation to specify a tensor program at the kernel, thread block, and thread levels of the GPU compute hierarchy, and uses a novel pruning technique based on abstraction to significantly reduce the search space Mirage needs to consider while providing a certain optimality guarantee. Mirage outperforms existing tensor program optimizers by up to 3.5$\times$ even for DNNs that are widely used and heavily optimized.